OpenStudy (anonymous):

@Somy How many liters of oxygen gas, at standard temperature and pressure, will react with 35.4 grams of calcium metal? Show all of the work used to solve this problem. 2 Ca + O2 yields 2 CaO For this question would I use PV = nRT?

OpenStudy (somy):

no this time you do not use it

OpenStudy (somy):

find moles of calcium metal

OpenStudy (anonymous):

0.8

OpenStudy (somy):

there are 2 Ca

OpenStudy (somy):

40.1* 2= your Mr for Ca

OpenStudy (anonymous):

Oops, right. So like 2.26

OpenStudy (somy):

Mr = 80.2 mass = 35.4

OpenStudy (somy):

mole= mass/Mr try again

OpenStudy (anonymous):

0.44

OpenStudy (somy):

thats right)

OpenStudy (somy):

2 Ca + O2 yields 2 CaO now ratio whats the ratio between Ca and O2?

OpenStudy (anonymous):

2:1?

OpenStudy (somy):

good so if 2- 0.44 1- ?

OpenStudy (anonymous):

.22

OpenStudy (somy):

thats right now u got moles of O2

OpenStudy (somy):

there is a formula connecting Volume , standard temperature and pressure value, and moles

OpenStudy (somy):

value of standard temperature and pressure is 22.4 if i remember correctly

OpenStudy (somy):

check out your notes

OpenStudy (anonymous):

Let me check real quick

OpenStudy (anonymous):

This is what I have

OpenStudy (somy):

oh its not about this the value is in Liters and it's known so u can straightaway use 22.4dm3

OpenStudy (anonymous):

OH, okay lol.

OpenStudy (somy):

\[Volume = mole \times 22.4 \] do this and you'll get your answer :)

OpenStudy (somy):

it's the formula for when you need to find volume or mole and you are given either one and the question is telling you that it's happening (reaction is happening) under standard temperature and pressure so do not mix it up :)

OpenStudy (anonymous):

V= .22 x 22.4?

OpenStudy (somy):

yes :)

OpenStudy (anonymous):

So it's 4.93

OpenStudy (somy):

that's right) and the unit is dm3 or Liters

OpenStudy (anonymous):

liters since it's asking for that right

OpenStudy (somy):

yup) so u don't need to convert anything :D

OpenStudy (anonymous):

Awesome! Can you help me with my last one?

OpenStudy (somy):

lets see it :)

OpenStudy (anonymous):

If 5.0 liters of carbon dioxide gas are produced by the reaction below at STP, how many liters of oxygen gas were used in the reaction? 2 CO (g) + O2 (g) yields 2 CO2 (g)

OpenStudy (somy):

use volume formula :D

OpenStudy (somy):

and find mole of CO2

OpenStudy (anonymous):

V = 12 x 22.4?

OpenStudy (anonymous):

nvm wait a sec lol

OpenStudy (somy):

u are given volume of CO2 and u know STP so find mole

OpenStudy (anonymous):

5.0 = .22 x 22.4 = 4.9

OpenStudy (somy):

omg lol 0.22???? it's from ur previous question

OpenStudy (somy):

Volume= mole * 22.4 mole= Volume/22.4 =5.0/22.4=

OpenStudy (anonymous):

LOL im all over the place

OpenStudy (anonymous):

Wait so Im confused now

OpenStudy (somy):

what u don't get?